# Incremental Joint Learning of Depth, Pose and Implicit Scene Representation on Monocular Camera in Large-scale Scenes

### v2

@@ III-C Incremental Implicit Scene Representation
where $\delta_{i}$ is the distance between sample points and $T_{i}$ represents the accumulated transmittance along the ray. Compared with existing scene representation methods, we have re-parameterized the scene representation for unbounded scenes similar to [47]. We use a contract function to map every point into a cubic space:
@@
Depth Guided Sampling  The estimated depth images provide valuable geometry information which can guide neural point sampling along a ray. We get $N_{strat}$ points for stratified sampling between the near and far planes. Then, $N_{surface}$ points are drawn from the Gaussian distribution determined by the depth prior. When the depth is not known or invalid, we use the rendered depth. Compared to the original methods, our approach allows for more efficient point sampling and enhances the scene representation capability of the network.
@@ III-D Incremental Joint Learning
We sample M pixels for the frame. Then, we normalize the depth:
@@
We use the progressive optimization scheme to improve the robustness and to satisfy large-scale scenes and long sequences. We start by using a small amount of frames (for example five frames) to initiate the operation. Subsequently, we progressively introduce new frames into the optimization process in a sequential manner.
TABLE III: The view synthesis results on Tanks and temples dataset. We report the average PSNR, SSIM, LPIPS results.
@@ IV Experiments
Baselines Methods and Public Datasets. The main baseline methods we compared are BARF [3], NoPe-NeRF [9], and LocalRF [10]. All of them are joint learning methods of camera poses and implicit scene representation. We also compare our method against SCNeRF [4], NeRFmm [38], NeRFfacto [48]. We evaluate our method on the Tanks and Temples dataset [8] and Static Hikes dataset [10]. The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes. We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 100 meters. We strictly follow the experimental settings of our baseline to partition the training set and the test set.
@@
Our Dataset. An in-house proprietary dataset was collected using an intelligent wheelchair to simulate the application of NeRF in elderly care settings, as shown in Fig. 5. The dataset includes a long-distance travel scene spanning over 35 meters and consists of over 1k images. Our intelligent wheelchair features two sensors: the Realsense D435i, an RGBD camera capturing only the RGB images, and the Robosense RS-Lidar-16, aimed at obtaining groundtruth poses with a certified performance at 2 cm level accuracy [49]. We evaluate the capabilities of our depth, pose estimation, and scene representation. We use this dataset to validate the performance of our method in real-world robot operations.
@@
Figure 5: The image above is a prototype of our robot platform, and below are the qualitative results of our method on our own real-world datasets with PSNR value. We compare the view synthesis results with NoPe-NeRF[9], LocalRF [10]. Our method performs better on real-world datasets.
@@
Implementation Details. We present our implementation details. We select Adam optimizer[50] ($\beta=(0.9,0.999)$) for scene representation and camera tracking optimization. The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth and $\lambda_{3}=1.0$ for pose estimation, $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow. All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU. Our initial learning rate are $5\times 10^{-3}$ for rotations and $5\times 10^{-4}$ for translations. The resolution of our local radiance fields is $64^{3}$. When the camera moves to the bound, we refine the local radiance fields and the camera pose for 600 iterations per frame and upsamle the resolution to $640^{3}$. After that, we allocate new local radiance fields and repeat this process until all frames are registered.
@@
TABLE IV: Novel view synthesis on Static Hikes dataset [10] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
@@ IV-A Experimental Results
Scene Reconstruction Results. We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate view synthesis results. In Table I, we present the scene reconstruction results on Tanks and Temples datasets. Our method outperforms all the baselines by a large margin. In Tab. III, we compare our method with other view synthesis methods with COLMAP pose, which shows the proves the advantages of our joint learning system in large-scale scenes. The scene reconstruction results on other datasets are shown in Table IV. Our method outperforms other methods in scene representation. The incremental implicit scene representation method significantly enhances the scene representation ability and performs well on large-scale and long-sequence datasets. The qualitative results are shown in Fig. 3,4,5. In Fig. 3, We present the view synthesis results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [8]. Our method achieves higher-quality image reconstruction results. In Fig. 4, these images correspond to different timesteps (90, 390, 900) from up to down. We place the PSNR values of each scene in the bottom right corner. In contrast, our incremental scene representation method achieves high-quality view synthesis results compared with NoPe-NeRF and LocalRF. In Fig. 5, we show the image reconstruction results on our own dataset. Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
@@
Figure 6: Qualitative results of depth estimation on Tanks and Temples datasets [8]. We present the depth estimation results of NoPe-NeRF[9], LocalRF [10], and our method. The region outlined on the image is marked in red and black to signify lower accuracy, in green to signify higher accuracy and groundtruth.
@@
Pose Estimation Results. The pose estimation results are shown in Table. II and Table. IV. We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics. Our pose estimation network performs better than other existing methods. The FBA method performs re-projection optimization in the high-dimensional feature level, which avoids various disturbances and noise present at the image level. The FBA network enhances the accuracy and robustness of the pose estimation in large-scale scenes. We also present the depth estimation results in Fig. 6. Compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks. Our estimated depth images are more accurate and sharper, as shown in green blocks.
@@
TABLE V: Ablation studies of different modules on Tanks and Temples datasets [8].
@@ IV-B Ablation Study
In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network. We show the ablation results in Table V. (a) we remove the depth estimation and depth prior. (b) is the network without incremental scene representation. (c) is the network without feature-metric bundle adjustment. (d) is our method using pose from COLMAP. In Table (a), we remove the depth loss and the depth network is not jointly optimized with other networks. The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation. (b) For the incremental scene representation, we replace it with global optimization. We can see that the incremental scene representation significantly improves the scene representation and pose estimation. We also present the render results between using a single global radiance field and using multiple radiance fields in Fig. 7. This indicates the importance of our incremental scene representation method. (c) We remove the feature-metric bundle adjustment and replace it with traditional pose optimization. We can see that this network improves the accuracy of the pose estimation results. (d) We use the camera pose from COLMAP and fix the camera pose throughout the entire optimization. This demonstrates that our joint optimization method outperforms the original method based on COLMAP.
@@ V Conclusion
In this paper, we propose a novel incremental joint learning network of depth, pose, and implicit scene representation. The proposed FBA network extracts multi-level image features and performs coarse-to-fine BA estimation for corresponding pixels and image pose, which can provide accurate pose estimation results both in indoor and outdoor scenes. An incremental scene representation method is designed for large-scale scenes and long-sequence videos. This proposed network can achieve accurate and robust scene representation for arbitrarily long-sequence scene reconstruction. Compared with other joint learning methods, our method achieves SOTA results in depth estimation, pose estimation, and implicit scene representation in large-scale indoor and outdoor scenes.

### v3

@@ III-C Incremental Implicit Scene Representation
scenes BARF[2] NoPe-NeRF [7] LocalRF [8] Flow-NeRF [53] RA-NeRF [54] Ours
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Tanks and Temples Ballroom 20.66 0.50 0.30 24.97 0.71 0.20 25.29 0.89 0.08 28.83 0.86 0.24 30.25 N/A 0.11 26.50 0.85 0.08
Barn 25.28 0.64 0.24 25.77 0.68 0.22 27.89 0.88 0.11 28.53 0.78 0.35 25.99 N/A 0.38 28.59 0.91 0.11
Church 23.17 0.62 0.27 23.60 0.67 0.23 28.07 0.91 0.07 28.27 0.83 0.28 32.05 N/A 0.08 28.82 0.92 0.08
Family 23.04 0.61 0.28 23.77 0.68 0.22 28.60 0.92 0.007 29.40 0.85 0.29 30.46 N/A 0.15 30.48 0.94 0.05
Francis 25.85 0.69 0.28 29.48 0.80 0.18 31.71 0.93 0.11 30.63 0.83 0.33 32.04 N/A 0.24 33.38 0.93 0.10
Horse 24.09 0.72 0.21 27.30 0.83 0.13 29.98 0.94 0.07 28.57 0.86 0.23 28.26 N/A 0.18 30.07 0.93 0.07
Ignatius 21.78 0.47 0.30 23.77 0.61 0.23 26.54 0.86 0.09 26.25 0.73 0.35 27.77 N/A 0.25 27.83 0.87 0.09
Museum 23.58 0.61 0.28 26.39 0.76 0.17 26.35 0.86 0.16 29.43 0.85 0.27 31.59 N/A 0.10 29.43 0.92 0.12
mean 23.42 0.61 0.27 26.34 0.74 0.19 27.87 0.91 0.09 28.73 0.82 0.29 29.80 N/A 0.18 29.84 0.90 0.09
Local Scene Representation Voxel grid-based architectures [55] are the mainstream in NeRF-based systems. However, they struggle with cubical memory growing and space complexity. In our local scene representation, we factorize the local radiance fields to compact components and reduce the space complexity from cubic ($\mathcal{O}(n^{3})$) to square ($\mathcal{O}(n)$). The 3D feature grid $\mathcal{G}$ can be factorized as with Vector-Matrix (VM) decomposition:
$\mathcal{G}=\sum_{r=1}^{R}\mathbf{v}_{r}^{X}\circ\mathbf{M}_{r}^{YZ}+\mathbf{v}_{r}^{Y}\circ\mathbf{M}_{r}^{XZ}+\mathbf{v}_{r}^{Z}\circ\mathbf{M}_{r}^{XY}$ (7)
where $\mathbf{v}_{r}^{X}$,$\mathbf{v}_{r}^{Y}$, $\mathbf{v}_{r}^{Z}$ are the vectors corresponding to XYZ axis. Our factorization-based model efficiently computes the feature vector for each voxel at a low cost, requiring only a single value per XYZ-mode vector or matrix factor. Additionally, we incorporate a tri-linear interpolation method to achieve a continuous fields. In our local neural radiance field, we can just interpolate the vector/matirx factor for the corresponding modes. Instead of recovering eight individual tensor elements for trilinear interpolation, we directly compute the interpolated value, significantly reducing computation and memory costs during runtime.
@@
where $\delta_{i}$ is the distance between sample points, $T_{i}$ represents the accumulated transmittance along the ray, and $N$ denotes the number of sample points along the ray. Compared with existing scene representation methods, we have re-parameterized the scene representation for unbounded scenes similar to [56]. We use a contract function to map every point into a cubic space:
@@
Depth Guided Sampling  The estimated depth images provide valuable geometry information which can guide neural point sampling along a ray. We get $N_{strat}$ points for stratified sampling between the near and far planes. Then, $N_{surface}$ points are drawn from the Gaussian distribution determined by the depth prior. When the depth is not known or invalid, we use the rendered depth. Compared to the original methods, our approach allows for more efficient point sampling and enhances the scene representation capability of the network. We illustrate the differences between our sampling method and original sampling method in Fig. 3.
@@ III-D Incremental Joint Learning
We sample N pixels for the frame. Then, we normalize the depth:
@@
We use the progressive optimization scheme to improve the robustness and to satisfy large-scale scenes and long sequences. This joint optimization framework simultaneously leverages 2D photometric consistency from RGB images and 3D structural information from depth, enforcing coherence between 2D appearance and 3D geometry. This synergy leads to enhanced accuracy in depth estimation, pose prediction, and overall scene reconstruction. We start by using a small amount of frames (for example five frames) to initiate the operation. Subsequently, we progressively introduce new frames into the optimization process in a sequential manner.
TABLE III: The depth estimation accuracy on the Tanks and Temples dataset [52] and indoor datasets [8], using Depth L1, RMSE, Absolute Relative Error (Abs Rel), and Square Relative Error (Sq Rel) as metrics.
Scenes NoPe-NeRF [7] LocalRF [8] Ours
[…]
Depth-L1 [cm] $\downarrow$ 159.57 176.79 158.77
RMSE [cm] $\downarrow$ 171.50 191.45 171.69
[…]
Depth-L1$\downarrow$ 187.27 153.88 136.2
[…]
TABLE IV: The scene reconstruction results on Tanks and temples dataset. We report the average PSNR, SSIM, LPIPS results.
@@ IV Experiments
Baselines Methods and Public Datasets. The main baseline methods we compared are BARF [2], NoPe-NeRF [7], LocalRF [8], Flow-NeRF [53], RA-NeRF [54]. All of them are joint learning methods of camera poses and implicit scene representation. We also compare our method against SCNeRF [3], NeRFmm [37], NeRFfacto [57]. We evaluate our method on the Tanks and Temples dataset [52] (from 4x4m to 8x8m) and Static Hikes dataset [8] (15x15m). The Tanks and Temples dataset features slow-moving image sequences spanning a few meters in both indoor and outdoor scenes. We also utilize the Static Hikes indoor dataset, which includes hand-held sequences with longer camera trajectories, testing scalability and pose estimation robustness over ranges typically exceeding 100 meters (30x30m). We strictly follow the experimental settings of our baseline to partition the training set and the test set.
@@
Our Dataset. An in-house proprietary dataset was collected using an intelligent wheelchair to simulate the application of NeRF in elderly care settings, as shown in Fig. 6 and Fig. 7. We collected datasets and conducted experiments in both laboratory environments and real hospital scenarios. The dataset includes a long-distance travel scene spanning over one hundred meters and consists of over 1k images. Our intelligent wheelchair features two sensors: the Realsense D435i, an RGBD camera capturing only the RGB images, and the Robosense RS-Lidar-16, aimed at obtaining groundtruth poses with a certified performance at 2 cm level accuracy [58]. We evaluate the capabilities of our depth, pose estimation, and scene representation. We use this dataset to validate the performance of our method in real-world robot operations.
@@
Figure 6: The image above is a prototype of our robot platform, and below are the qualitative results of our method on our own real-world datasets with PSNR value. We compare the scene reconstruction results with NoPe-NeRF[7], LocalRF [8]. Our method performs better on real-world datasets.
Figure 7: This image shows the reconstruction results of different methods in a real robot platform within a hospital scenario. We compare the scene reconstruction results with NoPe-NeRF[7], LocalRF [8]. Our approach adapts better to data from various scenes, achieving more accurate scene reconstruction and pose estimation.
@@
Implementation Details. We present our implementation details. We select Adam optimizer[59] ($\beta=(0.9,0.999)$) for scene representation and camera tracking optimization. The color loss weighting is $\lambda_{1}=0.25$, $\lambda_{2}=0.1$ for depth, $\lambda_{3}=1.0$ for pose estimation, and $\lambda_{4}=1.0$, $\lambda_{5}=1.0$ for optical flow. All training and evaluation experiments are conducted on a single NVIDIA RTX 3090 GPU. Our initial learning rate are $5\times 10^{-3}$ for rotations and $5\times 10^{-4}$ for translations. The resolution of our local radiance fields is $64^{3}$. When the camera moves to the bound, we refine the local radiance fields and the camera pose for 600 iterations per frame and upsamle the resolution to $640^{3}$. After that, we allocate new local radiance fields and repeat this process until all frames are registered.
@@
TABLE V: Novel view synthesis on Static Hikes dataset [8] and Proprietary dataset. We use PSNR, SSIM, and LPIPS as our metrics.
@@ IV-A Experimental Results
Scene Reconstruction Results. We use Peak Signal-to-noise Ratio (PSNR), Structural Similarity Index Measurement (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to evaluate scene reconstruction results. In Table I, we present the scene reconstruction results on Tanks and Temples datasets. Our method outperforms all the baselines by a large margin, thanks to our incremental scene representation method and efficient tri-plane representation. In Tab. IV, we compare our method with other scene reconstruction methods with/without COLMAP pose, which shows the proves the advantages of our joint learning system in large-scale scenes. The scene reconstruction results on Static Hikes and our own datasets are shown in Table V. Our method outperforms other methods in scene representation. The incremental implicit scene representation method significantly enhances the scene representation ability and performs well on large-scale and long-sequence datasets. The qualitative results are shown in Fig. 4,5,6. In Fig. 4, we present the scene reconstruction results on different sequences Ignatius, Horses, Church in Tanks and Temples dataset [52]. Our method achieves higher-quality image reconstruction results. In Fig. 5, these images correspond to different timesteps (90, 390, 900) from up to down. We place the PSNR values of each scene in the bottom right corner. In contrast, our incremental scene representation method achieves high-quality scene reconstruction results compared with NoPe-NeRF and LocalRF. In Fig. 6 and Fig. 7, we show the image reconstruction results in our own dataset. Compared with other joint learning methods, our method achieves the most high-fidelity novel view results in both outdoor and indoor scenes.
Figure 8: Visualization of pose estimation results on Tanks and Temples datasets [52]. We present the pose estimation results of BARF [2], NoPe-NeRF[7], LocalRF [8], and our method. We visualize the Absolute Trajectory Error ATE of different methods. The color bar on the right represents the ATE pose error.
@@
Figure 9: Qualitative depth estimation results on the family sequence from the Tanks and Temples dataset [52]. We present the depth estimation results of NoPe-NeRF[7], LocalRF [8], and our method.
Figure 10: Qualitative results of depth estimation on Indoor datasets [52]. We present the depth estimation results of NoPe-NeRF[7], LocalRF [8], and our method.
@@
Pose Estimation Results. The pose estimation results are shown in Table. II and Table. V. We use Absolute Trajectory Error (ATE), and Relative Pose Error (RPE) as the pose estimation metrics. Our pose estimation network performs better than other existing methods. We also visualize the ATE error of different methods in Tanks datasets. The FBA method performs re-projection optimization in the high-dimensional feature level, which avoids various disturbances and noise present at the image level. The FBA network enhances the accuracy and robustness of the pose estimation in large-scale scenes.
Depth Estimation Results. We present the depth estimation results in Fig. 9,10. In Fig. 9, compared to our method, a tree is missing from the LocalRF estimated depth, as shown in red blocks, and the depth at both bushes and trees in NoPe-NeRF is inaccurate, as shown in black blocks. Our estimated depth images are more accurate and sharper, as shown in green blocks. We present quantitative results in Tab. III. Our depth estimation performance is better than other NeRF-Based methods thanks to the joint learning methods.
@@
TABLE VI: Ablation studies of different modules on Tanks and Temples datasets [52].
@@ IV-B Ablation Study
In this section, we conduct sufficient ablation studies to verify the effectiveness of our designed network. We show the ablation results in Table VI. (a) we remove the depth estimation and depth prior. (b) is the network without incremental scene representation. (c) is the network without feature-metric bundle adjustment. (d) is our method using pose from COLMAP. In Table (a), we remove the depth loss and the depth network is not jointly optimized with other networks. The results indicate that the depth loss and joint optimization strategy are effective for scene representation and pose estimation. (b) For the incremental scene representation, we replace it with global optimization. We can see that the incremental scene representation significantly improves the scene representation and pose estimation. We also present the render results between using a single global radiance field and using multiple radiance fields in Fig. 11. This indicates the importance of our incremental scene representation method. (c) We remove the feature-metric bundle adjustment and replace it with traditional pose optimization. We can see that this network improves the accuracy of the pose estimation results. (d) We use the camera pose from COLMAP and fix the camera pose throughout the entire optimization. This demonstrates that our joint optimization method outperforms the original method based on COLMAP.
@@ V Conclusion
In this paper, we propose a novel incremental joint learning network of depth, pose, and implicit scene representation. The proposed FBA network extracts multi-level image features and performs coarse-to-fine bundle adjustment estimation for corresponding pixels and image pose, which can provide accurate pose estimation results both in indoor and outdoor scenes. An incremental scene representation method is designed for large-scale scenes and long-sequence videos. This proposed network can achieve accurate and robust scene representation for arbitrarily long-sequence scene reconstruction. Compared with other joint learning methods, our method achieves SOTA results in depth estimation, pose estimation, and implicit scene representation in large-scale indoor and outdoor scenes.
Future work Currently, our experiments are primarily conducted in relatively large-scale indoor environments. As future work, we plan to extend our approach to even larger, city-scale scenes. To ensure better global consistency in such expansive environments, it is also necessary to incorporate techniques such as loop closure detection into our framework.